# On the representation-free formalism in quantum theory

### v10

@@ II.3 Features of the bra-ket scheme
Further, one needs to consider the drawbacks of the above-described revised version of the bra-ket scheme. First, let us examine the conventional bra-ket expressions for matrix elements, such as $\langle u|O|v\rangle$. The reasoning on this subject in this paragraph is basically similar to the considerations by Gieres gie . The expressions $\langle u|O|v\rangle$ is introduced through the equalities (1). However, these equalities are not universally valid. Indeed, let $D(O)$ and $D(O^{\dagger})$ denote the domains of definition of the operators $O$ and $O^{\dagger}$. The hermiticity relation (4), on which the equalities (1) are based, applies only when $|v\rangle\in D(O)$ and $|u\rangle\in D(O^{\dagger})$ simultaneously. To provide a complete definition of the $O^{\dagger}$ operator, the domain $D(O^{\dagger})$ is to be specified. It is to be accepted that the (maximal) domain $D(O^{\dagger})$ consists of all the $u$ vectors for which the relation (4) takes place for any $v$ belonging to $D(O)$. Suppose that $u\notin D(O^{\dagger})$. In this case, the above definition of the covector $\langle u|O$ is unapplicable, and the matrix element $\bigl{(}\langle u|O\bigr{)}|v\rangle$ entering Eq. (1) does not exist. Suppose now that $v\notin D(O)$. Then the matrix element $\langle u|\bigl{(}O|v\rangle\bigr{)}$ that appears in Eq. (1) does not exist. In both cases, the expression $\langle u|O|v\rangle$ is undefined. Nevertheless, the matrix elements $(u,Ov)$ can exist even when $u\notin D(O^{\dagger})$ (assuming $D(O^{\dagger})\neq{\cal H}$) and the matrix elements $(O^{\dagger}u,v)$ can exist even when $v\notin D(O)$ (assuming $D(O)\neq{\cal H}$). In such cases, the well-defined matrix elements $(u,Ov)$ and $(O^{\dagger}u,v)$ cannot be represented by the conventional bra-ket expression $\langle u|O|v\rangle$.
@@ III Present formulation
For the quantum mechanical scalar product, we suggest the expression like $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ or, for instance, ${\bf r}/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/plm/$, or ${\bf p}_{1},\ldots,{\bf p}_{n}/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$. While in the bra-ket scalar product expression $\langle u|v\rangle$ the bra vector $\langle u|$ is a linear functional, in the present $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ expression $u$ is a vector. Thus, $u$ and $v$ belong to the same space. The $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ expression is an analog to the expression for the scalar product of vectors in finite-dimensional spaces, such as ${\bf u}\cdot{\bf v}$. (Of course, $\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u=(u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v)^{*}, vv$$\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v\geq 0$, and $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ is linear in $v$ and thus anti-inear in $u$.) The adjoint operator $O^{\dagger}$ of a linear operator $O$ is defined in the conventional way:
$O^{\dagger}u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}Ov$ (5)
where the domain of definition of $O^{\dagger}$ is indicated in relation to Eq. (4). (If $O$ is anti-linear then $O^{\dagger}u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=Ov\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u$.) Eq. (5) is simpler than Eq. (4). Unlike the bra-ket formalism, adjoint operators are not a component of the present scheme. For this reason and others, this scheme is less involved than the bra-ket one.
@@
In some (though not all) cases, it is convenient to omit the dot in the scalar products of the form $u/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O/v/$ writing them simply as /u/O/v//u/O/v/. (One has $u/O/v/^{*}=O/v/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/u/$.) Matrix elements of an operator in a basis are then represented like
@@
We introduce multiplication signs * between constituents of emerging expressions, such as scalar products or state vectors and scalar products. This appears as $u*v\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}w$, or $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v*u^{\prime}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v^{\prime}*w$. The multiplication signs provide clarity, and in particular, they are required for what follows.
One should distinguish between the expressions $cu\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ and $(cu)\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$, where $c$ is a constant. While the second expression represents the scalar product of the states $cu$ and $v$, the first one is interpreted as the product of the constant $c$ and the scalar product $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ (unlike expressions such as $Ou\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$). One has $(\sum_{i}c_{i}u_{i})\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=\sum_{i}(c_{i}u_{i})\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=\sum_{i}c_{i}^{*}u_{i}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$. If, for example, in the above relations the constants $c$ or $c_{i}$ themselves are scalar products then for clarity the replacement $c$ by cc* or $c_{i}$ by $c_{i}*$ and $c_{i}^{*}$ by $c_{i}^{*}*$ is to be done.
Further, as in the case of the bra-ket formalism, we propose an ’explicit’ expression for projection-type operators. In the present scheme, this expression is of the form $u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. It incorporates the multiplication sign between the vectors $u$ and $v$ while the dot now follows the vector $v$. The result of the action of the operator  $u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ on a state $w$ is then expressed as $u*v\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}w$. Here, $v\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}w$ represents the scalar product as defined above. Thus, the action of projection-type operators is self-explanatory. It follows from Eq. (5) that $(u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}})^{\dagger}=v*u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$.
Let us write down some simple relations employing the above expression for projection-type operators. If $\{/m/\}$ is an orthonormal basis, then the identity operator $I$ can be expressed as $I=\sum_{m}/m/*/m/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. (Of course, this expression is equivalent to the expansions of state vectors over the set /m//m/, such as $u=\sum_{m}/m/*c_{m}$, where $c_{m}=/m/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u$.) Similarly, if $\nu/$ is, e.g., a complete set of continuous observables then $I$ can be written as $\int d\nu/\nu/*/\nu/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$.
@@
$q/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\psi\equiv/q/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}I\psi=\int d\nu/q/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/\nu/*/\nu/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\psi.$
@@
Let $\{/m/\}$ and $\{/n/\}$ be orthonormal bases. One may formally represent any operator $O$ in terms of the projection-type operators $m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$:
$O=\sum_{m,n}\,/m/O/n/*/m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}.$
@@
In particular, one may write the identity operator $I$ as $\sum_{m,n}\,/m/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/n/*/m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$, which reduces to the expression for $I$ written above if the bases $\{/m/\}$ and $\{/n/\}$ coincide.
@@
If $O_{pt}$ is a projection-type operator then the product $O_{1}*O_{pt}*O_{2}$ of three operators can be formally represented as a single projection-type operator. Denote this latter operator as $\mu/*/\nu/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ and $O_{pt}$ as $m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. Then we have $\mu/=O_{1}/m/$ and $\nu/=O_{2}^{\dagger}/n/$.
@@
One may write the product of projection-type operators as such an operator times the product of scalar products, $\prod_{i=1}^{n}u_{i}*v_{i}\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}=\left(\prod_{i=1}^{n-1}v_{i}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u_{i+1}\right)u_{1}*v_{n}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. All this shows the utility of the explicit expression for projection-type operators.
@@
Up to this point, we formulated the present scheme as a one-space scheme. Let us incorporate the dual-space point of view in the present scheme. To this aim, we will reinterpret our notation. In the above scalar product expression $\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v, uu$ was considered a state vector, and the dot represented the scalar product sign. And now we will treat $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ as an indivisible symbol. Namely, we will consider $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ as the other designation for the linear functional $G_{u}$ in the relation $G_{u}(v)=u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$. That is, one will understand $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ as $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}(v)$ under this interpretation. Consequently, we will interpret $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ as a covector. This interpretation was proposed in the preliminary version of the present paper efr . [^7] In particular, it has been pointed out in efr (since the 2023 [v7] version) that an easy way to learn the bra-ket formalism is to proceed from this interpretation, as described below. A similar approach is also advocated in the recent paper lev . (The notation $Ou\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ has the same meaning as $(Ou)\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. In contrast, if $c$ is a constant then $cu\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ is interpreted as the product of the constant $c$ and the covector $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. If $c_{i}$ are constants and $u_{i}$ are state vectors, then we have $(\sum_{i}c_{i}u_{i})\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}=\sum_{i}c_{i}^{*}u_{i}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. If, for example, $c_{i}$ and $c_{i}^{*}$ are scalar products, then for clarity they should be replaced by $c_{i}*$ and $c_{i}^{*}*$.)
@@
In this way, one shifts to the dual-space description discussed in the previous section regarding the bra-ket formalism. According to this description, one deals with the spaces of state vectors and state covectors. The scalar product $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ becomes the ”inner product” of a state covector $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ and a state vector $v$. And in both one-space and dual-space versions of the present scheme, the projection-type operator $u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ is the ”outer product” of a state vector $u$ and a state covector $v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$.
@@
As mentioned in the previous section, the bra-ket expressions for covectors, such as $\langle\psi|O$ or $\langle\psi|O^{\dagger}$, entail potential ambiguities regarding the domain of definition. This drawback is absent here. Indeed, for instance, the covector $\langle\psi|O^{\dagger}$ is expressed as $O\psi\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ in the present formulation. From this expression, it is apparent that $\psi$ should belong to the domain of definition of the operator $O$.
(One can express a mapping $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\rightarrow O\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ of covectors in terms of an operator acting on the covectors $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. Let us denote this operator as $\bar{O}$ and position it to the right of covectors. Then, we have $O\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\equiv\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}{\bar{O}}$. Accordingly, scalar products of the form $O\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\chi$ can be rewritten as $(\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}{\bar{O}})\chi$. If the operator $O$ is linear, then the corresponding operator $\bar{O}$ is also linear. In terms of $\bar{O}$, the relation defining conjugate operators takes the form $(\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}{\bar{O}})\chi=\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O^{\dagger}\chi$. If this relation holds for any states $\psi$ and $\chi$, one may view the operator $\bar{O}$ as the operator $O^{\dagger}$ acting to the left, i.e., on covectors. Accordingly, in this case, one can denote $\bar{O}$ as $O^{\dagger}$, assuming that in the expression $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O^{\dagger}\chi$ the operator $O^{\dagger}$ may act both to the right and to the left. This applies, in particular, in the finite-dimensional case.)
@@
$|\psi\rangle\Leftrightarrow\psi,\qquad\langle\psi|\Leftrightarrow\psi\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}},\qquad\langle\psi|O\Leftrightarrow O^{\dagger}\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}.$
Further, when transitioning from the bra-ket formulation to the present formulation, the matrix element expression $\langle\psi|O|\chi\rangle$ should be replaced with either of the expressions $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O\chi$ and $O^{\dagger}\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\chi$, provided that each of the latter expressions is well-defined. And if one of these expressions is not well-defined, then the bra-ket expression $\langle\psi|O|\chi\rangle$ has no sense, as discussed in Sec. 2 C. An easy opportunity to master the widely used bra-ket formalism is to start with the one-space version of the present scheme and then move on to its dual-space version, utilizing the listed correspondence.

### v12

@@ II.3 Features of the bra-ket scheme
Further, one needs to consider the drawbacks of the above-described revised version of the bra-ket scheme. First, let us examine the conventional bra-ket expressions for matrix elements, such as $\langle u|O|v\rangle$. The reasoning on this subject in this paragraph is basically similar to the considerations by Gieres gie . The expressions $\langle u|O|v\rangle$ is introduced through the equalities (1). However, these equalities are not universally valid. Indeed, let $D(O)$ and $D(O^{\dagger})$ denote the domains of definition of the operators $O$ and $O^{\dagger}$. The hermiticity relation (4), on which the equalities (1) are based, applies only when $|v\rangle\in D(O)$ and $|u\rangle\in D(O^{\dagger})$ simultaneously. To provide a complete definition of the $O^{\dagger}$ operator, the domain $D(O^{\dagger})$ is to be specified. It is to be accepted that the (maximal) domain $D(O^{\dagger})$ consists of all the $u\equiv|u\rangle$ vectors for which the relation (4) takes place for any $v\equiv|v\rangle$ vector belonging to $D(O)$. Suppose that $u\notin D(O^{\dagger})$. In this case, the above definition of the covector $\langle u|O$ is unapplicable, and the matrix element $\bigl(\langle u|O\bigr)|v\rangle$ entering Eq. (1) does not exist. Suppose now that $v\notin D(O)$. Then the matrix element $\langle u|\bigl(O|v\rangle\bigr)$ that appears in Eq. (1) does not exist. In both cases, the expression $\langle u|O|v\rangle$ is undefined. Nevertheless, the matrix elements $(u,Ov)$ can exist even when $u\notin D(O^{\dagger})$ (assuming $D(O^{\dagger})\neq{\cal H}$) and the matrix elements $(O^{\dagger}u,v)$ can exist even when $v\notin D(O)$ (assuming $D(O)\neq{\cal H}$). In such cases, the well-defined matrix elements $(u,Ov)$ and $(O^{\dagger}u,v)$ cannot be represented by the conventional bra-ket expression $\langle u|O|v\rangle$.
@@ III Present formulation
For the quantum mechanical scalar product, we suggest the expression like $u\boldsymbol{\cdot}v$ or, for instance, ${\bf r}/\boldsymbol{\cdot}/plm/$, or ${\bf p}_{1},\ldots,{\bf p}_{n}/\boldsymbol{\cdot}v$. While in the bra-ket scalar product expression $\langle u|v\rangle$ the bra vector $\langle u|$ is a linear functional, in the present $u\boldsymbol{\cdot}v$ expression $u$ is a vector. Thus, $u$ and $v$ belong to the same space. The $u\boldsymbol{\cdot}v$ expression is an analog to the expression for the scalar product of vectors in finite-dimensional spaces, such as ${\bf u}\cdot{\bf v}$. (Of course, $\boldsymbol{\cdot}u=(u\boldsymbol{\cdot}v)^{*}, vv$$\boldsymbol{\cdot}v\geq 0$, and $u\boldsymbol{\cdot}v$ is linear in $v$ and thus anti-inear in $u$.) The adjoint operator $O^{\dagger}$ of a linear operator $O$ is defined in the conventional way:
$O^{\dagger}u\boldsymbol{\cdot}v=u\boldsymbol{\cdot}Ov$ (5)
where the domain of definition of $O^{\dagger}$ is indicated in relation to Eq. (4). (If $O$ is anti-linear then $O^{\dagger}u\boldsymbol{\cdot}v=Ov\boldsymbol{\cdot}u$.) Eq. (5) is simpler than Eq. (4). Unlike the bra-ket formalism, adjoint operators are not a component of the present scheme. For this reason and others, this scheme is less involved than the bra-ket one.
@@
In some (though not all) cases, it is convenient to omit the dot in the scalar products of the form $u/\boldsymbol{\cdot}O/v/$ writing them simply as /u/O/v//u/O/v/. (One has $u/O/v/^{*}=O/v/\boldsymbol{\cdot}/u/$.) Matrix elements of an operator in a basis are then represented like
@@
We introduce multiplication signs * between constituents of emerging expressions, such as scalar products or state vectors and scalar products. This appears as $u*v\boldsymbol{\cdot}w$, or $u\boldsymbol{\cdot}v*u^{\prime}\boldsymbol{\cdot}v^{\prime}*w$. The multiplication signs provide clarity, and in particular, they are required for what follows.
One should distinguish between the expressions $cu\boldsymbol{\cdot}v$ and $(cu)\boldsymbol{\cdot}v$, where $c$ is a constant. While the second expression represents the scalar product of the states $cu$ and $v$, the first one is interpreted as the product of the constant $c$ and the scalar product $u\boldsymbol{\cdot}v$ (unlike expressions such as $Ou\boldsymbol{\cdot}v$). One has $(\sum_{i}c_{i}u_{i})\boldsymbol{\cdot}v=\sum_{i}(c_{i}u_{i})\boldsymbol{\cdot}v=\sum_{i}c_{i}^{*}u_{i}\boldsymbol{\cdot}v$. If, for example, in the above relations the constants $c$ or $c_{i}$ themselves are scalar products then for clarity the replacement $c$ by cc* or $c_{i}$ by $c_{i}*$ and $c_{i}^{*}$ by $c_{i}^{*}*$ is to be done.
Further, as in the case of the bra-ket formalism, we propose an ’explicit’ expression for projection-type operators. In the present scheme, this expression is of the form $u*v\,\boldsymbol{\cdot}$. It incorporates the multiplication sign between the vectors $u$ and $v$ while the dot now follows the vector $v$. The result of the action of the operator  $u*v\,\boldsymbol{\cdot}$ on a state $w$ is then expressed as $u*v\boldsymbol{\cdot}w$. Here, $v\boldsymbol{\cdot}w$ represents the scalar product as defined above. Thus, the action of projection-type operators is self-explanatory. It follows from Eq. (5) that $(u*v\,\boldsymbol{\cdot})^{\dagger}=v*u\,\boldsymbol{\cdot}$.
Let us write down some simple relations employing the above expression for projection-type operators. If $\{/m/\}$ is an orthonormal basis, then the identity operator $I$ can be expressed as $I=\sum_{m}/m/*/m/\boldsymbol{\cdot}$. (Of course, this expression is equivalent to the expansions of state vectors over the set /m//m/, such as $u=\sum_{m}/m/*c_{m}$, where $c_{m}=/m/\boldsymbol{\cdot}u$.) Similarly, if $\nu/$ is, e.g., a complete set of continuous observables then $I$ can be written as $\int d\nu/\nu/*/\nu/\boldsymbol{\cdot}$.
@@
$q/\boldsymbol{\cdot}\psi\equiv/q/\boldsymbol{\cdot}I\psi=\int d\nu/q/\boldsymbol{\cdot}/\nu/*/\nu/\boldsymbol{\cdot}\psi.$
@@
Let $\{/m/\}$ and $\{/n/\}$ be orthonormal bases. One may formally represent any operator $O$ in terms of the projection-type operators $m/*/n/\boldsymbol{\cdot}$:
$O=\sum_{m,n}\,/m/*/m/\boldsymbol{\cdot}O/n/*/n/\boldsymbol{\cdot}\equiv/m/*/m/O/n/*/n/\boldsymbol{\cdot}\equiv/m/O/n/*/m/*/n/\boldsymbol{\cdot}.$
@@
In particular, one may write the identity operator $I$ as $\sum_{m,n}\,/m/\boldsymbol{\cdot}/n/*/m/*/n/\boldsymbol{\cdot}$, which reduces to the expression for $I$ written above if the bases $\{/m/\}$ and $\{/n/\}$ coincide.
@@
If $O_{pt}$ is a projection-type operator then the product $O_{1}*O_{pt}*O_{2}$ of three operators can be formally represented as a single projection-type operator. Denote this latter operator as $\mu/*/\nu/\boldsymbol{\cdot}$ and $O_{pt}$ as $m/*/n/\boldsymbol{\cdot}$. Then we have $\mu/=O_{1}/m/$ and $\nu/=O_{2}^{\dagger}/n/$.
@@
One may write the product of projection-type operators as such an operator times the product of scalar products, $\prod_{i=1}^{n}u_{i}*v_{i}\,\boldsymbol{\cdot}=\left(\prod_{i=1}^{n-1}v_{i}\boldsymbol{\cdot}u_{i+1}\right)u_{1}*v_{n}\boldsymbol{\cdot}$. All this shows the utility of the explicit expression for projection-type operators.
@@
Up to this point, we formulated the present scheme as a one-space scheme. Let us incorporate the dual-space point of view in the present scheme. To this aim, we will reinterpret our notation. In the above scalar product expression $\boldsymbol{\cdot}v, uu$ was considered a state vector, and the dot represented the scalar product sign. And now we will treat $u\,\boldsymbol{\cdot}$ as an indivisible symbol. Namely, we will consider $u\,\boldsymbol{\cdot}$ as the other designation for the linear functional $G_{u}$ in the relation $G_{u}(v)=u\boldsymbol{\cdot}v$. That is, one will understand $u\boldsymbol{\cdot}v$ as $u\boldsymbol{\cdot}(v)$ under this interpretation. Consequently, we will interpret $u\,\boldsymbol{\cdot}$ as a covector. This interpretation was proposed in the preliminary version of the present paper efr . [^7] In particular, it has been pointed out in efr (since the 2023 [v7] version) that an easy way to learn the bra-ket formalism is to proceed from this interpretation, as described below. A similar approach is also advocated in the recent paper lev . (The notation $Ou\,\boldsymbol{\cdot}$ has the same meaning as $(Ou)\,\boldsymbol{\cdot}$. In contrast, if $c$ is a constant then $cu\boldsymbol{\cdot}$ is interpreted as the product of the constant $c$ and the covector $u\boldsymbol{\cdot}$. If $c_{i}$ are constants and $u_{i}$ are state vectors, then we have $(\sum_{i}c_{i}u_{i})\boldsymbol{\cdot}=\sum_{i}c_{i}^{*}u_{i}\boldsymbol{\cdot}$. If, for example, $c_{i}$ and $c_{i}^{*}$ are scalar products, then for clarity they should be replaced by $c_{i}*$ and $c_{i}^{*}*$.)
@@
In this way, one shifts to the dual-space description discussed in the previous section regarding the bra-ket formalism. According to this description, one deals with the spaces of state vectors and state covectors. The scalar product $u\boldsymbol{\cdot}v$ becomes the ”inner product” of a state covector $u\,\boldsymbol{\cdot}$ and a state vector $v$. And in both one-space and dual-space versions of the present scheme, the projection-type operator $u*v\,\boldsymbol{\cdot}$ is the ”outer product” of a state vector $u$ and a state covector $v\,\boldsymbol{\cdot}$.
@@
As mentioned in the previous section, the bra-ket expressions for covectors, such as $\langle\psi|O$ or $\langle\psi|O^{\dagger}$, entail potential ambiguities regarding the domain of definition. This drawback is absent here. Indeed, for instance, the covector $\langle\psi|O^{\dagger}$ is expressed as $O\psi\,\boldsymbol{\cdot}$ in the present formulation. From this expression, it is apparent that $\psi$ should belong to the domain of definition of the operator $O$.
(One can express a mapping $\psi\boldsymbol{\cdot}\rightarrow O\psi\boldsymbol{\cdot}$ of covectors in terms of an operator acting on the covectors $\psi\boldsymbol{\cdot}$. Let us denote this operator as $\bar{O}$ and position it to the right of covectors. Then, we have $O\psi\boldsymbol{\cdot}\equiv\psi\boldsymbol{\cdot}{\bar{O}}$. Accordingly, scalar products of the form $O\psi\boldsymbol{\cdot}\chi$ can be rewritten as $(\psi\boldsymbol{\cdot}{\bar{O}})\chi$. If the operator $O$ is linear, then the corresponding operator $\bar{O}$ is also linear. In terms of $\bar{O}$, the relation defining conjugate operators takes the form $(\psi\boldsymbol{\cdot}{\bar{O}})\chi=\psi\boldsymbol{\cdot}O^{\dagger}\chi$. If this relation holds for any states $\psi$ and $\chi$, one may view the operator $\bar{O}$ as the operator $O^{\dagger}$ acting to the left, i.e., on covectors. Accordingly, in this case, one can denote $\bar{O}$ as $O^{\dagger}$, assuming that in the expression $\psi\boldsymbol{\cdot}O^{\dagger}\chi$ the operator $O^{\dagger}$ may act both to the right and to the left. This applies, in particular, in the finite-dimensional case.)
@@
$|\psi\rangle\Leftrightarrow\psi,\qquad\langle\psi|\Leftrightarrow\psi\,\boldsymbol{\cdot},\qquad\langle\psi|O\Leftrightarrow O^{\dagger}\psi\boldsymbol{\cdot}.$
Further, when transitioning from the bra-ket formulation to the present formulation, the matrix element expression $\langle\psi|O|\chi\rangle$ should be replaced with either of the expressions $\psi\boldsymbol{\cdot}O\chi$ and $O^{\dagger}\psi\boldsymbol{\cdot}\chi$, provided that each of the latter expressions is well-defined. And if one of these expressions is not well-defined, then the bra-ket expression $\langle\psi|O|\chi\rangle$ has no sense, as discussed in Sec. 2 C. An easy opportunity to master the widely used bra-ket formalism is to start with the one-space version of the present scheme and then move on to its dual-space version, utilizing the listed correspondence.